OpenStudy (anonymous):

The perimeter of a retangle is 34ft. the lenght is 3ft longer than the width. Find the demensions. Write a system of linear equation and solvethe resulting system. Let x be the length and y be the width. write the first equation

OpenStudy (anonymous):

I disagree with this answer. the dimensions will be 7 & 10. 4x=28

OpenStudy (anonymous):

if the rectangle measures 34, you have to subtract the extra 3 from the long sides and turn it into a square, this gives you 7 per side, now add the 3 back to your two longer sides. you will have to do your own linear equation, because mine would be wrong, 2x+[(2y)+6]=34 2x+2y=34-6 4x7=28 do you see what I mean? you have to do the linear equation

OpenStudy (anonymous):

so what is the next step do I divide

OpenStudy (anonymous):

2x=14 1x=7 x+3=y y=10 do have the first person's response? i don't know how to write it out so that your teacher will accept it.